Operations Report – February 26th

Crew 206 Operations Report 26 – FEB – 2019

SOL: 2

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Generator (hours run): Turned off at 0800 (24th Feb), Turned on at 1800 (24th Feb)

Solar— SOC % – Turned off Gen (24th Feb) 82%, Turned on Gen (24th Feb) 64%

Diesel Reading – ~ 75%

Propane Reading – ~ 23%

Ethanol Free Gasoline – 7 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – around 93%, ~ 509 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 285,1 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01415823 units

Toilet tank emptied: No

Deimos rover used: No, still not functional
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: Not Applicable
Beginning charge: Not Applicable
Ending charge: Not Applicable
Currently charging: Not Applicable

Opportunity rover used: Used
Hours: 51.7 hours
Beginning charge: 100%
Ending charge: 95%
Currently charging: Yes

Curiosity rover used: Used
Hours: 75.9 hours
Beginning charge: 100%
Ending charge: 93%
Currently charging: Yes

Spirit rover used: Not Used
Hours: 78.3 hours
Beginning charge: 100%
Ending charge: 100%
Currently charging: YES

Notes on rovers: Charging socket has been fixed, everything is currently fine.

ATV’s Used: N/A

Reason for use: Not used
Oil Added? N/A
ATV Fuel Used: N/A
# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Not used

Crew Car used and why, where? Off Campus

General notes and comments:

Summary of internet: Lost Internet access once in the day. Came back after no more than 10 minutes. Probably the same issue as yesterday.

Summary of suits and radios: Checked all the suits batteries. They are all fine and charging. Except for suit #11 on which the charging socket is disconnected from the battery. Will fix it ASAP. Checked suit #1 (had fogging issues yesterday) fans : they are working fine. Maybe the battery wasn’t charged enough for EVA #1.

Summary of Hab operations: Added some tape to the inside locker for engineering airlock. The door now closes more precisely.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAM operations:

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing more to report.

Copyright © The Mars Society. All rights reserved. | Main Site